need the answer as soon as possible with the steps

You borrow $20,000 from a bank to be repaid in monthly installments for 3 years at 9% interest compounded monthly. What is the interest portion of the 8th payment?

Group of answer choices

A/$88

B/$124

C/$84

D/$150
Answer #1

Given that loan amount is 20000, effective rate of interest is 9%/12 = 0.75% per month, total payments made = 3*12 = 36, we find that monthly payment is A = 20000(A/P, 0.75%, 36)

A = 20000*(0.0075*1.0075^36)/(1.0075^36-1)

= 635.99

Interest portion of 8th payment = 635.99*(P/A, 0.75%, 36 – 8 + 1)*0.75%

= 635.99*(1.0075^29-1)/(0.0075*1.0075^29)*0.0075

= $123.90

Select option B